# TopoBDA: Towards Bezier Deformable Attention for Road Topology Understanding

### v2

@@ 5 Conclusion
Experimental evaluations demonstrate that TopoBDA achieves state-of-the-art results across both subsets of the OpenLane-V2 dataset, utilizing the version 1.1 metric baseline. Specifically, TopoBDA surpasses existing methodologies with a DETl score of 38.9 and an OLS score of 51.7 in Subset-A, and a DETl score of 45.1 and an OLS score of 54.3 in Subset-B. The integration of multi-modal data significantly enhances performance: fusion of camera and lidar data elevates the OLS score in Subset-A from 51.7 to 56.4, and in Subset-B from 54.3 to 61.7. The subsequent incorporation of SDMap alongside camera and lidar sensors further augments the OLS score in Subset-A from 56.4 to 58.4. These findings underscore the efficacy of TopoBDA in comprehending road topology and highlight the substantial performance improvements afforded by sensor fusion. Additionally, TopoBDA achieves superior results on the OpenLane-V1 benchmark for 3D lane detection, with F1-Scores of 63.9 for the 1.5m distance and 57.9 for the 0.5m distance.
@@
In conclusion, this work aims to fill gaps in the literature, moving towards a comprehensive HDMap element prediction. TopoBDA represents a significant advancement in road topology understanding and 3D lane detection for autonomous driving. By leveraging Bezier deformable attention, instance mask formulation, multi-modal fusion, and an auxiliary one-to-many set prediction loss strategy, TopoBDA achieves superior performance in centerline detection and topology reasoning. This approach not only enhances computational efficiency but also sets a new benchmark in the field, highlighting its potential for practical applications in autonomous driving systems.

### v3

@@ 5 Conclusion
Experimental evaluations demonstrate that TopoBDA achieves state-of-the-art performance across both subsets of the OpenLane-V2 dataset, using the version 1.1 metric baseline. Specifically, TopoBDA surpasses existing methods with a DETl score of 38.9 and an OLS score of 51.7 in Subset-A, and a DETl score of 45.1 and an OLS score of 54.3 in Subset-B. The integration of multi-modal data significantly boosts performance: fusing camera and LiDAR data increases the OLS score in Subset-A from 51.7 to 56.4, and in Subset-B from 54.3 to 61.7. Further incorporating SDMap alongside camera and LiDAR sensors raises the OLS score in Subset-A to 58.4. These results underscore the effectiveness of TopoBDA in road topology comprehension and highlight the substantial benefits of multi-modal fusion.
Additionally, TopoBDA achieves superior results on the OpenLane-V1 benchmark for 3D lane detection, with F1-scores of 63.9 at a 1.5m distance and 57.9 at a 0.5m distance.
@@
This work contributes toward closing existing gaps in HDMap element prediction, offering a unified framework for road topology understanding and 3D lane detection in autonomous driving. By leveraging Bezier deformable attention, instance mask formulation, multi-modal fusion, and an auxiliary one-to-many set prediction loss strategy, TopoBDA delivers high accuracy in centerline detection and topological reasoning. The approach not only improves computational efficiency but also sets a new benchmark in the field, highlighting its potential for practical applications in autonomous driving systems. The contributions of the TopoBDA study are also demonstrated in Supplementary Table S9 in Section S.8.
The analysis of modality fusion in TopoBDA represents an important step toward comprehensive road topology understanding. Notably, this work introduces the integration of LiDAR data into road topology modeling for the first time. However, due to the high cost of LiDAR sensors, the advantages of this modality are primarily realized in offline settings, such as automated HD map extraction or ground-truth labeling of static elements. In contrast, the combination of SD maps and camera inputs within TopoBDA has also demonstrated significant benefits, indicating that an online deployment scenario on edge devices remains feasible.
Although BDA demonstrates superior efficiency compared to alternative attention mechanisms, the base TopoBDA architecture remains impractical for deployment on edge devices due to its computational complexity, particularly the number of encoder and decoder layers. The original architecture comprises six encoder layers and ten decoder layers, which is computationally demanding even for online execution on high-performance servers. Consequently, the selection of encoder and decoder layers should be tailored to the resource constraints of the target edge device. For the edge device scenario, to illustrate the impact of efficient multi-scale implementations and varying layer configurations, this study provides comparative analyses in Table S2 (Supplementary Section S.6.2) and in Table S4 and Table S5 (Supplementary Section S.6.3).
Despite its strengths, TopoBDA has certain limitations due to its Bezier keypoint-dependent structure. While it performs well in detecting structured elements such as centerlines and lane dividers, it may face challenges in drivable area (road boundary) prediction. These regions often exhibit complex geometries with protrusions and indentations, which are difficult to represent using a fixed number of Bezier control points. Accurately modeling complex road geometries may necessitate a higher number of control points; even the 8-control-point design of TopoBDA might be insufficient in certain cases, as evidenced by BeMapNet [qiao_end--end_2023]. BeMapNet addresses this limitation by employing piecewise Bézier curves for road boundary representation, suggesting that similar adaptations could be required for TopoBDA to achieve robust modeling of intricate road layouts.
Moreover, the architecture imposes constraints on feature dimensionality: the number of channels in query features must be divisible by the number of Bezier control points, which correspond to the number of heads in the cross-attention (Bezier Deformable Attention). Consequently, adjusting the number of control points may necessitate changes in feature dimensions, potentially complicating design flexibility. A detailed analysis of these constraints is provided in Supplementary Section S.6.5.